The world producer price for baseball is $24 per dozen, and almost all of them are produced


outside the United States. Suppose the U.S. demand curve and supply curves are:


QD = 100 000 – 2 000 P


QS = - 10 000 + 1 000 P


P : price per dozen


Q: dozens


1) Before a tariff is imposed, what is the U.S. equilibrium price? Domestic consumption?


Domestic production? imports?


2) Congress has decided to help the baseballmanufacturing industry by imposing a tariff of $ 6 per dozen. What are the new equilibrium price, domestic consumption, domestic production, and imports?


3) What are the losses to U.S. consumers, gains to U.S. producers, and deadweight loss? (Hint: The area of a trapezoid is [height] [ base 1 + base 2 ] / 2 where base 1 and base 2 are the parallel sides of the trapezoid.)


1
Expert's answer
2021-12-05T19:39:49-0500

Task #273282


(1). Before a tariff is imposed, what is the U.S. equilibrium price? Domestic consumption? Domestic production? imports?


Solution:


The domestic consumption equilibrium is:

"QD = 100 000 \u2013 2 000 P"

"P =\$24"

"QD = 100 000 \u2013 2 000 (24)= 52,000"


The domestic production equilibrium is : 


"QS = - 10 000 + 1 000 P"

"P=\$24"

"QS = - 10 000 + 1 000 (24)= 14,000"


Imports :


"I= 52,000-14,000=38,000"  

"I=38,000"  


(2) . Congress has decided to help the baseball manufacturing industry by imposing a tariff of $ 6 per dozen. What are the new equilibrium price, domestic consumption, domestic production, and imports?


Solution;


The new equilibrium price after the tariff is:


"P= 24 + 6=\$ 30"  


The domestic consumption equilibrium is:


"QD = 100 000 \u2013 2 000 P"

"P=\$30"

"QD = 100 000 \u2013 2 000 (30)= 40,000"


The domestic production equilibrium is :


"QS = - 10 000 + 1 000 P"

"P=\$30"

"QS = - 10 000 + 1 000 (30) = 20,000"


Imports fall from:


"I= 40,000-20,000= 20,000"  

"I= 20,000"


(3). What are the losses to U.S. consumers, gains to U.S. producers, and dead weight loss? (Hint: The area of a trapezoid is [height] [ base 1 + base 2 ] / 2 where base 1 and base 2 are the parallel sides of the trapezoid.)


Solution:


The loss in consumer surplus is the area of the trapezoid:


"\$6(\frac {40,000+52,000}{2}) =\$276,000"


The gain in producer surplus is the are of the trapezoid:


"\$6(\frac{20,000+14000}{2})= \$102,000"


Revenue earned by the government form the tariff:


"\$6(40,000-20000)=\$120,000"


Dead weight loss:


"\$276,000-\$102,000-\$120,000= \$54,000"